# ProGeny I: a new simple stellar population generator and impact of isochrones / stellar atmospheres / initial mass functions

### v2

@@ Abstract
In this work we introduce ProGeny, a new stellar population library (SPL) software package written in R. This release encapsulates the core software to generate simple/single stellar populations (SSPs); the various data inputs required (in particular isochrones and stellar atmospheres); example scripts to generate the SSPs; and a number of pre-generated static SSP available for immediate use. The most novel feature of ProGeny is its ability to produce SSPs with evolving initial mass functions (IMFs), allowing functional dependencies on stellar age or metallicity. We perform both internal comparisons (within the ProGeny SPL) and external comparisons (with other public SSPs) and tests. The main conclusion is that the choice of isochrone has significantly more impact on the predicted spectra than the choice of stellar atmospheres and/or IMF (comparing Chabrier and Kroupa variants). A number of limiting uncertainties and corrections for star formation rates and stellar masses are also presented.
@@
pubyear: 2024pagerange: ProGeny I: a new simple stellar population generator and impact of isochrones / stellar atmospheres / initial mass functionsB
@@ 1 Introduction
The experience of using ProSpect for analysis has motivated a deeper investigation into the precise role SSPs play when analysing data. While ProSpect does come with a number of available SSPs, their precise production varies in detail and it is not possible to fully control and understand the causes of any inference variations. To combat this, and in parallel provide more flexibility to the types of SSP available for use in ProSpect and any similar SED generation or fitting software, we have produced the new open source (LGPL-3) R [^1] R Core Team (2023) package ProGeny that we present in this work. It offers highly flexible production of SSPs via the variation of isochrones, stellar atmospheres and IMFs. Perhaps most novel, compared to any current literature SSPs, ProGeny provides the means to generate evolving IMF SSPs, either as a function of age or metallicity (see Marks et al., 2012; Martín-Navarro et al., 2015).
@@
The paper provides a brief overview of various aspects of SSP generation, but we note that more detailed references are readily available in the literature. We highlight the review of Conroy (2013), and the earlier work of Martins & Coelho (2007), which was the first thorough investigation into the role of stellar atmospheres in synthetic spectral modelling of galaxies. In this work we discuss in detail the available isochrones, spectral atmospheres and IMFs provided with this initial release of ProGeny.222v0.4.0 is the exact version used to create outputs for this paper and the associated Bellstedt & Robotham in prep. companion paper).
@@
While some fiducial SSPs outputs are generated in association with this paper (and are analysed and discussed in latter sections), detailed examples are also provided and documented so users can produce their own SSPs as required in under a minute of computational time. It is also relatively easy for advanced users to supplement the provided isochrones, stellar atmospheres and IMFs with their own, as long as they follow the required catalogue and FITS file formats. The package requires no compilation and has a small number of dependencies, so many common barriers to entry for experimentation have been removed. We note this paper is joined by a companion paper (Bellstedt & Robotham in prep.) that investigates in depth the impact that switching SPLs and ProGeny SSPs has when fitting galaxy SEDs.
@@
In this work we discuss the main aspects of generative stellar population synthesis modelling (Section 2); the specific version of the ProGeny SPL released in this work (Section 3); a comparison between different ingredients in ProGeny and other literature SSPs (Section 4); some overall conclusions of the work (Section 5); and a summary of the software, script and data released to coincide with this paper (‘Data and Software Availability’). Because of the large number of software packages and astronomy acronyms used in this paper, a compact glossary is included in Appendix A. Where relevant we assume a H0 = 67.8 (kms/s)/Mpc, $\Omega_{M}$ = 0.301 and $\Omega_{\Lambda}=0.699$ cosmology (Planck Collaboration et al., 2020).
@@ 2 Stellar Population Synthesis Methodology
Stellar population synthesis (SPS) is generally used as the encompassing term for the steps required to make a synthetic spectrum of a distribution of stars (e.g. Conroy, 2013). A specific approach to doing this is what we consider to be a stellar population library (SPL), and a specific combination of parameters from an SPL is what generates an SSP (in our nomenclature).
ProGeny embodies two core concepts: a flexible software methodology for defining isochrones, stellar atmospheres, and interpolation preferences; and an SPL of static SSPs333where in our usage an SPL is a family of related SSPs. We encourage inquisitive users to explore and modify default options to create customised SSPs suited to their needs, particularly the initial mass function (IMF). Users can also supplement the provided isochrones and stellar atmospheres with new or alternative versions. As ProGeny evolves, we will expand base data options, creating new varieties while maintaining the original reference library for posterity.
@@ 2.1 Isochrone Format
Isochrones are the standard means for conveying how stars evolve over time. Typically, they begin at the zero-age main-sequence (ZAM) point, marking the onset of star formation, and trace the evolution of a stellar population to a target age (and/or end of life). Some isochrones extend their coverage to pre-main-sequence (PMS) phases, modelling the evolution of low-mass objects that never reach hydrogen fusion (e.g., brown and red dwarfs). Others incorporate special treatment for more exotic periods of stellar evolution (e.g. horizontal branch and asymptotic giant branch stars), and/or the evolution of remnant material (e.g. black holes, white dwarfs and neutron stars).
The key information that isochrones must convey are the age, mass, luminosity (and potentially the effective physical size), temperature and specific surface gravity of the relevant stars. Ideally, this is also done for a range a metallicity (Z) and $\alpha$ abundances, since both of these properties have a material impact on the evolution of stars. In a perfect world this information is provided in infinitesimally small increments of all the most interesting parameters for the purposes of constructing an SSP: age, mass and Z. However, pragmatically choices have to be made about the coarseness of the sampling in all dimensions either due to data and modelling limitations, or due to the lack of resulting spectral variability making higher resolution sampling unnecessary. ProGeny only requires regular binning in terms of age and metallicity, i.e. the stellar masses for each of these subsets can vary. In this way, the target isochrone adjusts the discrete masses to focus on those undergoing the most dramatic phases of stellar evolution, such as the highly luminous horizontal branch or asymptotic giant branch phases. This is how all the included MIST, PARSEC, and BaSTI isochrones operate (discussed in detail in Section 3.1).
@@
Table 1: Example of the minimal isochrone tabular information required for ProGeny. Columns are named as expected. ‘logZ’ is the logarithmic metallicity relative to solar; ‘logAge’ is the logarithmic age in years; ‘Mini’ is the initial formation mass of the star; ‘Mass’ is the current mass of the star (which will always be the same or lower than ‘Mini’ due to various mass loss events); ‘Lum’ is the luminosity in solar units; ‘Teff’ is the effective temperature of the photosphere in kelvin; ‘logG’ is the logarithmic specific gravity in CGS units ($\text{cm}/\text{s}^{2}$); ‘label’ is the stellar phase label (this is optional, but it can be used to match stellar atmospheres more accurately). This specific extract is from the 1,494,453 row MIST isochrone included with ProGeny.
@@
2.2 Stellar Atmosphere Spectral Format
When making an SPL, the equally important complement to the choice of isochrones is the selection of stellar atmospheres, i.e. the predicted spectral outputs of stars for specific combinations of stellar evolutionary parameters. The main properties that determine the precise spectrum of a stellar photosphere are (in roughly descending order of importance) the: effective surface temperature (Teff), the effective surface gravity (g), the stellar metallicity (Z) and the alpha abundance ($\alpha$). The ideal version of ProGeny would have stellar atmospheres generated for the exact combination of these variables required by the isochrone at all ages, however this is never possible in practice because the isochrones track the evolution of stars (and therefore these key parameters) much more finely than could ever be generated by a stellar atmosphere library. The reason for this depends on the source of the stellar atmospheres: for theoretical templates the limitation is really one of computation time and data volume (Hauschildt et al., 1999), and for observed templates the limitation is the difficulty observing enough stars with a broad high resolution spectrograph (e.g. X-Shooter on the VLT; Vernet et al., 2011). The pragmatic solution is then to assemble stellar atmospheres as densely as is feasible, and at the desired resolution of the target SSP.
The fiducial set of stellar atmospheres that are provided with ProGeny are discussed in more detail in Section 3.2. All stellar atmospheres share a common FITS file format that provides ProGeny with the critical information needed for properly constructing an SSP. Should users wish to supplement ProGeny with additional stellar atmospheres, or simply replace the fiducial libraries with different options, this is the multi-extensions FITS format that must be replicated:
@@ 2.2 Stellar Atmosphere Spectral Format
  • info: FITS table extension containing columns of Teff (Kelvin); logG (log10 surface gravity, g/[cm/s2]); logZ (log10 metallicity, Z) logA (log10 alpha abundance, $\alpha$). The rows must match 1–1 with the rows of the fourth extension spec.
@@
  • spec: FITS matrix of the stellar atmospheres bolometrically normalised to 1. The rows should correspond to the properties of second extension table info, and the columns are the relevant spectra that match 1–1 with the first extension vector wave.
@@ 2.3 Grid Interpolation Scheme
The common problem all SPL software that generate SSPs have to solve is a mechanism to match the very fine resolution predictions of stellar properties generated by isochrones onto the necessarily coarser grids of available stellar atmosphere spectra. High-quality isochrones capture small shifts in the effective temperate (Teff) and the surface gravity (logG), and ideally map out variations in metallicity (logZ) and other properties such as $\alpha$ abundances and stellar rotation.
@@
Being pragmatic, the best solution is to map out a very broad range of these properties with a few different libraries (since different libraries tend to focus on different properties, and none cover all the parameter space generated by modern isochrones) and then use interpolation methods to approximate specific solutions that fall between available grid positions. Previous efforts have generally focussed on direct spectral interpolation (where each wavelength element is treated as a multi-dimensional interpolation problem, e.g. Verro et al., 2022) or global interpolation (where entire spectra are weighted depending on their ‘distance’ to the objective stellar properties, e.g. Bruzual & Charlot, 2003). Which method is used is normally informed by the spectral data in hand: empirically focussed libraries tend to be so sparse that direct spectral interpolation is critical, but theoretically focussed libraries tend to have much denser grids of spectra (especially in the crucial Teff and logG dimensions) so global interpolation is preferable. ProGeny has been predominantly developed using theoretical stellar atmospheric spectra as inputs, so we have focussed on global interpolation strategies.
@@
Having chosen a global interpolation scheme, there are a number of additional subjective decisions to be made. How do we decide ‘distance’ in such multi-dimensional space, and what type of penalty do we want to apply? It is clear that spectra which are natively closer to the target properties produced for a particular star in our isochrone would always be preferred, but preferring spectra that are, e.g.  100K closer in Teff but 0.1 dex further away in logG is not as obvious.
@@
We carried out tests for how best we could replicate missing spectra. The optimal strategy in general is to define ‘distance’ relative to the average gridding separation, in either linear of log units as appropriate. Distance in each parameter dimension ($d_{i}$) is then defined as relative to the mean grid distance ($\mu_{\Delta\nu_{i}}$) in this dimension $i$, where the upper Euclidian radius (or $L^{2}$ norm) we search over is 2 (by default) in this unit-less re-mapping of the parameter space. We then weight the spectra by the inverse of the re-scaled distance to the power 2 (by default), so if a library atmosphere is twice as close in this re-scaled parameter space it will contribute four times as much flux. For a particular spectrum and a particular parameter dimension with an absolute target to grid separation of $\delta\nu_{i}$, the weight will look like:
@@
We can choose to limit the number of neighbouring spectra we allow to contribute (by default we limit this to 8, corresponding to cubic vertices in three-dimensional space) and enforce the sum of weights to equal 1 via a final renormalisation after all nearest matches have been combined.
When producing a spectrum for a target age stellar population we can combine all the weights for common spectral interpolation matches. This means we can stack spectra much more efficiently than simply adding every occurrence with a small weight, instead adding a particular stellar atmosphere spectrum once with all the weight terms summed accordingly. This interpolation method uses Arya et al. (1998) Approximate Nearest Neighbours (ANN) C++ library, so it is highly efficient for interpolating spectral grids even onto very large isochrone tables — 1.5m MIST isochrone parameters takes about 2 seconds to map onto the 7.5k stellar atmosphere grid values of the default Husser et al. (2013) library discussed in detail below. The parameter interpolation only needs to be carried out once per generation of a new SSP, and in fact only needs to be redone if the search radius or weight power terms are being modified (i.e. a change of IMF would not require the regeneration of interpolation weights).

### v3

# ProGeny I: a new simple stellar population spectra generator and impact of isochrones / stellar spectra / initial mass functionsD

@@ Abstract
In this work we introduce ProGeny, a new stellar population library (SPL) software package written in R. This release encapsulates the core open source software (github.com/asgr/ProGeny) to generate simple/single stellar populations and their associated spectra (SSPs); the various data inputs required (in particular isochrones and stellar spectra); example scripts to generate the SSPs; and a number of pre-generated static SSPs available for immediate use. The most novel feature of ProGeny is its ability to produce SSPs with evolving initial mass functions (IMFs), allowing functional dependencies on stellar age or metallicity. We perform both internal comparisons (within the ProGeny SPL) and external comparisons (to SSPs associated with other public SPLs) and tests. The main conclusion is that the choice of isochrone has significantly more impact on the predicted SSP spectra than the choice of stellar spectra and/or IMF (comparing Chabrier and Kroupa variants). A number of limiting uncertainties and corrections for star formation rates and stellar masses are also presented.
@@ 1 Introduction
The experience of using ProSpect for analysis has motivated a deeper investigation into the precise role SSPs play when analysing data. While ProSpect does come with a number of available SSPs, their precise production varies in detail and it is not possible to fully control and understand the causes of any inference variations. To combat this, and in parallel provide more flexibility to the types of SSP available for use in ProSpect and any similar SED generation or fitting software, we have produced the new open source (LGPL-3) R [^1] R Core Team (2023) package ProGeny that we present in this work. It offers highly flexible production of SSPs via the variation of isochrones, stellar spectra [^2] For clarity, this refers to the spectrum of a stellar photosphere, be it empirical or theoretical. In this paper when we refer to ‘stellar spectra’ or similar, this is what is meant. and IMFs. Perhaps most novel, compared to any current literature SSPs, ProGeny provides the means to generate evolving IMF SSPs, either as a function of age or metallicity (see Marks et al., 2012; Martín-Navarro et al., 2015).
@@
The paper provides a brief overview of various aspects of SSP spectra generation, but we note that more detailed references are readily available in the literature. We highlight the review of Conroy (2013), and the earlier work of Charlot et al. (1996); Maraston (2005); Martins & Coelho (2007); Percival & Salaris (2009); Conroy et al. (2009). The literature consensus has long been that within the bounds of reasonable ranges, the choice of isochrones has the most impact on the computed spectral outputs of SSPs, outranking the impact of stellar spectra and IMF. The stability and sensitivity of these ingredients also has a knock-on in terms of the accuracy specific quantities can be measured to. Generally stellar masses can be measured with a good degree of confidence (0.1–0.2 dex error being common, see Taylor et al., 2011; Robotham et al., 2020), but associated properties like age and metallicity are increasingly difficult to infer (Charlot et al., 1996). In this work we discuss in detail the available isochrones, stellar spectra and IMFs provided with this initial release of ProGeny.333v0.4.0 is the exact version used to create outputs for this paper and the associated Bellstedt & Robotham in prep. companion paper. A companion paper (ProGeny II; Bellstedt & Robotham in prep.) explores the impact of all these ingredients in depth, investigates how SSP choices affect the interpretation of observational data, and contains most of the ProGeny relevant analysis regarding SSP modelling error etc.
@@
While some fiducial SSPs outputs are generated in association with this paper (and are analysed and discussed in latter sections), detailed examples are also provided and documented so users can produce their own SSPs as required in under a minute of computational time. It is also relatively easy for advanced users to supplement the provided isochrones, stellar spectra and IMFs with their own, as long as they follow the required catalogue and FITS file formats. The package requires no compilation and has a small number of dependencies, so many common barriers to entry for experimentation have been removed. We note this paper is joined by a companion paper (Bellstedt & Robotham in prep.) that investigates in depth the impact that switching SPLs and ProGeny SSPs has when fitting galaxy SEDs.
@@
In this work we discuss the main aspects of generative stellar population synthesis modelling for single stars (i.e. we neglect binary treatment in this initial release) (Section 2); the specific version of the ProGeny SPL released in this work (Section 3); a comparison between different ingredients in ProGeny and other literature SSPs (Section 4); some overall conclusions of the work (Section 5); and a summary of the software, script and data released to coincide with this paper (‘Data and Software Availability’). Because of the large number of software packages and astronomy acronyms used in this paper, a compact glossary is included in Appendix D. Where relevant we assume a H0 = 67.8 (kms/s)/Mpc, $\Omega_{M}$ = 0.301 and $\Omega_{\Lambda}=0.699$ cosmology (Planck Collaboration et al., 2020). We note that when we refer to Z we implicitly mean log10(Fe/H) + 12 $\sim$ 7.46 and log10(O/H) + 12 $\sim$ 8.69 (as per Asplund et al., 2009). All isochrones and stellar spectra used are broadly consistent with this definition of Z, and therefore implied metallicity (always in units of Z/Zin ProGeny) can be scaled as desired.
@@ 2 Stellar Population Synthesis Methodology
Stellar population synthesis (SPS) is generally used as the encompassing term for the steps required to make a synthetic spectrum of a distribution of stars (e.g. Conroy, 2013). A specific approach to doing this is what we consider to be a stellar population library (SPL), and a specific combination of parameters from an SPL is what generates spectra for a SSP (in our nomenclature). For simplicity, when we refer to a SSP in this paper we are also referring to associated spectra generated (this is consistent with the nomenclature of Conroy, 2013).
ProGeny embodies two core concepts: a flexible software methodology for defining isochrones, stellar spectra, and interpolation preferences; and an SPL of semi-static SSPs444where in our usage an SPL is a family of related SSPs which will be improved and expanded over time. We encourage inquisitive users to explore and modify default options to create customised SSPs suited to their needs, particularly the initial mass function (IMF). Users can also supplement the provided isochrones and stellar spectra with new or alternative versions. As ProGeny evolves, we will expand base data options, creating new varieties while maintaining the original reference library for posterity.
@@ 2.1 Isochrone Format
Isochrones are the standard means for conveying how stars evolve over time. Typically, they begin at the zero-age main-sequence (ZAM) point, marking the onset of star formation, and trace the evolution of a stellar population to a target age (and/or end of life). Some isochrones extend their coverage to pre-main-sequence (PMS) phases, modelling the evolution of low-mass objects that never reach hydrogen fusion (e.g. brown and red dwarfs). Others incorporate special treatment for more exotic periods of stellar evolution (e.g. horizontal branch and asymptotic giant branch stars), and/or the evolution of remnant material (e.g. black holes, white dwarfs and neutron stars).
The key information that isochrones must convey are the age, mass, luminosity (and potentially the effective physical size), temperature and surface gravity of the relevant stars. Ideally, this is also done for a range a metallicity (Z) and $\alpha$/Fe abundances, since both of these properties have a material impact on the evolution of stars. In a perfect world this information is provided in infinitesimally small increments of all the most interesting parameters for the purposes of constructing an SSP: age, mass and Z. However, pragmatically choices have to be made about the coarseness of the sampling in all dimensions either due to data and modelling limitations, or due to the lack of resulting spectral variability making higher resolution sampling unnecessary. ProGeny only requires regular binning in terms of age and metallicity, i.e. the stellar masses for each of these subsets can vary. In this way, the target isochrone adjusts the discrete masses to focus on those undergoing the most dramatic phases of stellar evolution, such as the highly luminous horizontal branch or asymptotic giant branch phases. This is how all the included MIST, PARSEC, and BaSTI isochrones operate (discussed in detail in Section 3.1).
@@
Table 1: Example of the minimal isochrone tabular information required for ProGeny. Columns are named as expected. ‘logZ’ is the logarithmic metallicity relative to solar; ‘logAge’ is the logarithmic age in years; ‘Mini’ is the initial formation mass of the star; ‘Mass’ is the current mass of the star (which will always be the same or lower than ‘Mini’ due to various mass loss events); ‘Lum’ is the luminosity in solar units; ‘Teff’ is the effective temperature of the photosphere in Kelvin; ‘logG’ is the logarithmic surface gravity in CGS units ($\text{cm}/\text{s}^{2}$); ‘label’ is the stellar phase label (this is optional, but it can be used to match stellar spectra more accurately). This specific extract is from the 1,494,453 row MIST isochrone included with ProGeny.
@@
2.2 Stellar Photosphere Spectrum Format
When making an SPL, the equally important complement to the choice of isochrones is the selection of stellar spectra, i.e. the predicted spectral outputs of stars for specific combinations of stellar evolutionary parameters. The main properties that determine the precise spectrum of a stellar photosphere are (in roughly descending order of importance) the: surface temperature, the surface gravity, the stellar metallicity, and the alpha abundance ($\alpha$/Fe). The ideal version of ProGeny would have stellar spectra generated for the exact combination of these variables required by the isochrone at all ages, however this is never possible in practice because the isochrones track the evolution of stars (and therefore these key parameters) much more finely than could ever be generated by a stellar spectrum library. The reason for this depends on the source of the stellar spectra: for theoretical templates the limitation is really one of computation time and data volume (Hauschildt et al., 1999), and for observed templates the limitation is the difficulty observing enough stars with a broad high resolution spectrograph (e.g. X-Shooter on the VLT; Vernet et al., 2011). The pragmatic solution is then to assemble stellar spectra as densely as is feasible, and at the desired resolution of the target SSP.
The fiducial set of stellar spectra that are provided with ProGeny are discussed in more detail in Section 3.2. All stellar spectra share a common FITS file format that provides ProGeny with the critical information needed for properly constructing an SSP. Should users wish to supplement ProGeny with additional stellar spectra, or simply replace the fiducial libraries with different options, this is the multi-extensions FITS format that must be replicated:
@@ 2.2 Stellar Photosphere Spectrum Format
  • info: FITS table extension containing columns of Teff (Kelvin); logG (log10 surface gravity, g/[cm/s2]); logZ (log10 metallicity, Z) logA (log10 alpha abundance, $\alpha$/Fe). The rows must match 1–1 with the rows of the fourth extension spec.
@@
  • spec: FITS matrix of the stellar spectra bolometrically normalised to 1. The rows should correspond to the properties of second extension table info, and the columns are the relevant spectra that match 1–1 with the first extension vector wave.
@@ 2.3 Grid Interpolation Scheme
The common problem all SPL software that generate SSPs have to solve is a mechanism to match the very fine resolution predictions of stellar properties generated by isochrones onto the necessarily coarser grids of available stellar spectra. High-quality isochrones capture small shifts in the temperate and the surface gravity, and ideally map out variations in metallicity and other properties such as $\alpha$/Fe abundances and stellar rotation.
@@
Being pragmatic, the best solution is to map out a very broad range of these properties with a few different libraries (since different libraries tend to focus on different properties, and none cover all the parameter space generated by modern isochrones) and then use interpolation methods to approximate specific solutions that fall between available grid positions. Previous efforts have generally focussed on direct spectral interpolation (where each wavelength element is treated as a multi-dimensional interpolation problem, e.g. Verro et al., 2022) or global interpolation (where entire spectra are weighted depending on their ‘distance’ to the objective stellar properties, e.g. Bruzual & Charlot, 2003). Which method is used is normally informed by the spectral data in hand: empirically focussed libraries tend to be so sparse that direct spectral interpolation is critical, but theoretically focussed libraries tend to have much denser grids of spectra (especially in the crucial temperature and surface gravity dimensions) so global interpolation is preferable. ProGeny has been predominantly developed using theoretical stellar atmospheric spectra as inputs, so we have focussed on global interpolation strategies.
@@
Having chosen a global interpolation scheme, there are a number of additional subjective decisions to be made. How do we decide ‘distance’ in such multi-dimensional space, and what type of penalty do we want to apply? It is clear that spectra which are natively closer to the target properties produced for a particular star in our isochrone would always be preferred, but preferring spectra that are, e.g.  100 K closer in Teff but 0.1 dex further away in surface gravity is not as obvious.
@@
We carried out tests for how best we could replicate missing spectra. The optimal strategy in general is to define ‘distance’ relative to the average gridding separation, in either linear of log units as appropriate. Distance in each parameter dimension ($d_{i}$) is then defined as relative to the mean grid distance ($\mu_{\Delta\nu_{i}}$) in this dimension $i$, where the upper Euclidian radius (or $L^{2}$ norm) we search over is 2 (by default) in this unit-less re-mapping of the parameter space. We then weight the spectra by the inverse of the re-scaled distance to the power 2 (by default), so if a library stellar spectrum is twice as close in this re-scaled parameter space it will contribute four times as much flux. For a particular spectrum and a particular parameter dimension with an absolute target to grid separation of $\delta\nu_{i}$, the weight will look like:
@@
We can choose to limit the number of neighbouring spectra we allow to contribute (by default we limit this to 8, corresponding to cubic vertices in three-dimensional space) and enforce the sum of weights to equal 1 via a final renormalisation after all nearest matches have been combined. When determining the most appropriate spectrum to use ProGeny can optionally use information from the isochrone regarding the stellar evolutionary phase. This means that only a star flagged as being in the AGB phase would be assigned an AGB spectrum etc. This option is recommended, on by default, and used for all work in this paper.
When producing a spectrum for a target age stellar population we can combine all the weights for common spectral interpolation matches. This means we can stack spectra much more efficiently than simply adding every occurrence with a small weight, instead adding a particular stellar spectrum once with all the weight terms summed accordingly. This interpolation method uses Arya et al. (1998) Approximate Nearest Neighbours (ANN) C++ library, so it is highly efficient for interpolating spectral grids even onto very large isochrone tables — 1.5m MIST isochrone parameters takes about 2 seconds to map onto the 7.5k stellar spectrum grid values of the default Husser et al. (2013) library discussed in detail below. The parameter interpolation only needs to be carried out once per generation of a new SSP, and in fact only needs to be redone if the search radius or weight power terms are being modified (i.e. a change of IMF would not require the regeneration of interpolation weights).